# Multiple Emulsions (W/O/W) for Confined Precipitation of Drug Nanoparticles

### v1

@@ Abstract
Multiple emulsions offer a compelling route to confine nucleation and growth during drug precipitation, yet their practical use is frequently limited by kinetic fragility and sensitivity to formulation and processing conditions. Here, we develop an ultrasound-assisted, two-stage emulsification strategy to generate water-in-oil-in-water (W/O/W) multiple emulsions with sufficient stability to function as templates for forming drug-rich submicron particulates. We first establish an operating window using simple W/O emulsions, showing that increased Tween 80 concentration and intensified sonication (higher amplitude and larger probe) yield smaller droplets and reduced coarsening tendencies. Using this window, W/O/W emulsions are formulated and systematically screened via surfactant pairing across ionic, non-ionic, and polymeric stabilizers. Ionic–non-ionic combinations provide the most favorable droplet-size control, with CTAB–Tween 80 emerging as a practically robust formulation. Cyclohexane was selected as a reproducible platform oil for downstream precipitation using the lead CTAB–Tween 80 formulation. Finally, curcumin-loaded W/O/W constructs generate curcumin-rich submicron particulates, supporting multiple emulsions as experimentally accessible microstructured environments for particle engineering of poorly soluble drugs.

### v2

# arXiv:2601.06693v2[cond-mat.soft] 02 Mar 2026

@@ Abstract
Multiple emulsions offer a route to confine nucleation and growth during drug precipitation, yet practical use is often limited by kinetic fragility and sensitivity to formulation and processing conditions. Here, we develop an ultrasound-assisted, two-stage emulsification strategy to generate water-in-oil-in-water (W/O/W) multiple emulsions with sufficient stability to enable formation of drug-rich submicron particulates. We first establish an operating window using simple W/O emulsions, showing that increased Tween 80 concentration and intensified sonication (higher amplitude and larger probe) yield smaller droplets and reduced coarsening tendencies. Using this window, W/O/W emulsions are formulated and systematically screened via surfactant pairing across ionic, non-ionic, and polymeric stabilizers. Ionic–non-ionic combinations provide the most favorable droplet-size control, with CTAB–Tween 80 emerging as a practically robust formulation. Cyclohexane is identified as a reproducible platform oil for downstream precipitation using the lead CTAB–Tween 80 formulation. Finally, curcumin-loaded W/O/W constructs generate curcumin-rich submicron particulates, supporting multiple emulsions as experimentally accessible microstructured environments for particle engineering of poorly soluble drugs.
multiple emulsions; ultrasound emulsification; surfactant pairing; submicron particulates; curcumin precipitation
Impact Statement
Access to effective medicines is frequently limited by poor aqueous solubility of modern drug molecules, which can reduce and destabilize oral absorption and force higher doses, increasing cost and side effects. This work advances a practical, low-temperature route to improve drug dissolution performance by engineering drug-rich submicron particulates using water–oil–water (W/O/W) multiple emulsions as confined “microenvironments” for precipitation. We establish a two-stage, ultrasound-assisted emulsification workflow that decouples high-shear formation of a fine primary emulsion from low-shear transfer to the multiple-emulsion architecture, improving short-term kinetic robustness and preserving compartmentalization needed for controlled particle formation. Systematic screening identifies formulation and processing levers—surfactant concentration, ultrasound amplitude, probe geometry, and surfactant pairing—that jointly govern droplet stability and the ability to maintain confined domains. Using curcumin as a model poorly water-soluble compound, we demonstrate proof-of-concept formation of curcumin-rich submicron particulates within stabilized W/O/W constructs.
The approach is relevant to public health and sustainable manufacturing because it can reduce reliance on high-temperature or high-pressure particle-engineering routes, and it emphasizes scalable, experimentally accessible control variables (energy input, formulation composition). These features support translation toward cost-effective formulations that improve oral delivery of poorly soluble drugs, aligning with SDG 3 (Good Health and Well-Being) and SDG 9 (Industry, Innovation, and Infrastructure).